OpenStudy (anonymous):

Solve the equation... (x-1)(5x+2)=0

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

x=1 and x=-2/5

OpenStudy (anonymous):

That is what I thought, but how did you come up with -2/5?

OpenStudy (zehanz):

You have an equation of the form \(a \cdot b=0\). To solve it, break it up into: \(a=0\) or \(b=0\). Then solve each of these separately.

OpenStudy (anonymous):

2x+5=0 hence 2x=-5 and x=-5/2

OpenStudy (anonymous):

its other way

OpenStudy (anonymous):

5x+2=0 5x=-2 x=-2/5

OpenStudy (anonymous):

Wow...that was pretty simple....thanks for your help!!

OpenStudy (anonymous):

welcome...